# No “Zero-Shot” Without Exponential Data: Pretraining Concept Frequency Determines Multimodal Model Performance

### v2

@@ 8 Conclusions and Open Problems
In this work, we delved into the five pretraining datasets of 34 multimodal vision-language models, analyzing the distribution and composition of concepts within, generating over 300GB of data artifacts that we publicly release. Our findings reveal that across concepts, significant improvements in zero-shot performance require exponentially more data, following a log-linear scaling trend. This pattern persists despite controlling for similarities between pretraining and downstream datasets or even when testing models on entirely synthetic data distributions. Further, all tested models consistently underperformed on the “Let it Wag!” dataset, which we systematically constructed from our findings to test for long-tail concepts. This underlines a critical reassessment of what “zero-shot” generalization entails for multimodal models, highlighting the limitations in their current generalization capabilities. We highlight a few exciting avenues for future research to bridge these gaps or obtain further insights:
Understanding Image-Text Misalignments. One can explore the origins of misalignments between images and texts, such as the limitations of exact matching for concept identification in captions, inaccuracies from the RAM++ tagging model, or captions that are either too noisy or irrelevant.
Investigating Compositional Generalization. The term “zero-shot generalization” often refers to models’ ability for compositional generalization, meaning the ability to understand new combinations of concepts not previously encountered. This is distinct from traditional zero-shot learning and presents an intriguing, yet unresolved challenge: analyzing compositional generalization from a data-centric perspective.
@@ Acknowledgements
The authors would like to thank (in alphabetic order): Jonathan Roberts, Karsten Roth, Mehdi Cherti, Prasanna Mayilvahanan, Shyamgopal Karthik and Thao Nguyen for helpful feedback and providing access to various resources throughout the project. YS would like to thank Nicholas Carlini, Daphne Ippolito, Katherine Lee, Matthew Jagielski, and Milad Nasr. AP is funded by Meta AI Grant No. DFR05540. VU and YS thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS). VU also thanks the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. PT thanks the Royal Academy of Engineering for their support. AB acknowledges the Amazon Research Award. SA is supported by a Newton Trust Grant. MB acknowledges financial support via the Open Philanthropy Foundation funded by the Good Ventures Foundation. This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP4, project number: 276693517 and the UKRI grant: Turing AI Fellowship EP/W002981/1. MB is a member of the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.
@@ Appendix C Concept Frequency is Predictive of Performance for T2I Models
We extend the results from Fig. 3 with Figs. 11, 12, 13, 14 and 15. As with Fig. 3, due to the high concept frequency, the scaling trend is weaker. Furthermore, we do see inconsistency in the trends for the human-rated scores retrieved from HEIM [71], hence we perform a small scale human evaluation to check them.
@@
We assessed image-text alignment with a human study with 6 participants, where each participant was assigned 72 samples; for consistency, of the 360 total samples, we ensured 10% were assigned to 3 participants. Provided with a reference image, the participants were asked if the sample accurately depicts the prompt. Three choices were provided: “Yes” (score=1.), “Somewhat” (score=0.5), and “No” (score=0.). Accuracy was computed by averaging the scores.
@@ Appendix D Concept Frequency is Predictive of Performance across Concepts only from Image and Text Domains
In all the main performance-frequency plots we have presented until now, the concept frequencies were estimated using the intersection of the image-frequencies and the text-frequencies. Here, we showcase results with using them independently in Figs. 18 and 17 respectively. We note that both independent searching methods showcase log-linear trends as before confirming our main result. We observe that the strong log-linear trend between concept frequency and zero-shot performance robustly holds across concepts derived from image and text domains independently as well.
@@
Appendix E Experimental Details
E.1 Setup of Mayilvahanan et al. [79]
@@ Appendix E Experimental Details
LAION-200M is a dataset obtained by deduplicating LAION-400M by pruning exact duplicates, near duplicates, and semantically similar samples within LAION-400M [10]. The control pretraining set is created by pruning 50 million highly similar samples from LAION in the order of decreasing perceptual similarity to datapoints in ImageNet-val set. We use the 150M pretraining set for obtaining the concept distribution. We evaluate the performance of a ViT-B/32 CLIP model trained on this dataset on our downstream tasks, and present our analysis on those tasks.
E.2 Let It Wag!: Test Set Curation
@@
To ensure our datasets are thoroughly cleaned and diverse, we follow a meticulous process:
@@
1. Diverse Sourcing: We gather images from three different online sources—Flickr, DuckDuckGo, and Bing Search—to maximize the variety of our dataset, while retaining very easy-to-classify images222we use the image sourcing pipeline of C2C [90].
@@
2. Temporal Filtering: We applied a filter to only retrieve images after January 2023 to minimize overlap with images used in the pre-training of Vision-Language Models (VLMs). Note this helps mitigate but does not ensure the overlap problem is resolved.
@@ Appendix F Why and How Do We Use RAM++?
F.1 Why RAM++ and not CLIP or open-vocabulary detectors?
@@
F.2 How: Optimal RAM++ threshold for calculating concept frequencies
@@
We ablate the choice of the threshold we use for assigning concepts to images using the RAM++ model. For the given set of concepts, RAM++ provides a probability value (by taking a sigmoid over raw logits) for each concept’s existence in a particular image. To tag an image as containing a particular concept, we have to set a threshold deciding this assignnment. We test over three thresholds: {0.5, 0.6, 0.7}, showcasing quantitative and qualitative results for all thresholds in Figs. 21 and 20.
@@
We observe best frequency estimation results using the highest frequency of 0.7. This is due to the high precision afforded by this threshold, leading to us counting only the “most aligned images” per concept as hits. With lower thresholds (0.5, 0.6), we note that noisier images that do not align well with the concept can be counted as hits, leading to degraded precision and thereby poorer frequency estimation. Hence, we use 0.7 as the threshold for all our main results.
@@
Figure 20: Qualitative Results with different RAM++ thresholds. We show qualitative examples across three different thresholds: {0.5, 0.6, 0.7} for estimating concept frequency using the RAM++ model. We note that the significantly better concepts identified by the higher threshold (0.7) compared to the lower thresholds (0.5, 0.7). The images are sourced from the CC-3M dataset.
@@
Figure 21: Effect of different thresholds for determining concept frequency using RAM++. We test three different thresholds: {0.5, 0.6, 0.7} for estimating concept frequency using the RAM++ model. We note that the correlations are significantly stronger with a threshold of 0.7—this is justified by the higher precision of image sample hits at a higher threshold (0.7). Comparatively, lower thresholds (0.5, 0.7) lead to noisier images being counted as hits, hence reducing the hit precision for determining frequency. ** indicates that the result is significant ($p<0.05$ with two-tailed t-test.), and thus we show pearson correlation ($\rho$) too.
Appendix G Details about Misalignment Degree Results

### v3

@@ 8 Conclusion
In this work, we studied 5 pretraining datasets of 34 multimodal models, analyzing the distribution and composition of concepts within them, generating over 300GB of data artifacts that we publicly release. Our findings reveal that across concepts, significant improvements in zero-shot performance require exponentially more data, following a sample-inefficient log-linear scaling trend. This pattern persists despite controlling for similarities between pretraining and downstream datasets or even when testing models on entirely synthetic data distributions. Further, all tested models consistently underperformed on the “Let it Wag!” dataset, which we systematically constructed from our findings to test for long-tail concepts. This underlines a critical reassessment of what “zero-shot” generalization entails for multimodal models, highlighting the limitations in their current generalization capabilities.
@@ Acknowledgements
The authors would like to thank (in alphabetic order of first name): Gyungin Shin, Jonathan Roberts, Karsten Roth, Mehdi Cherti, Prasanna Mayilvahanan, Shyamgopal Karthik, Thao Nguyen and Vlad Bogolin for helpful feedback and providing access to various resources throughout the project. YS would like to thank Wieland Brendel, Nicholas Carlini, Daphne Ippolito, Katherine Lee, Matthew Jagielski, and Milad Nasr. AP is funded by Meta AI Grant No. DFR05540. VU and YS thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS). VU also thanks the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. VU was supported by a Google PhD Fellowship in Machine Intelligence. PT thanks the Royal Academy of Engineering for their support. AB acknowledges the funding from the KAUST Office of Sponsored Research (OSR-CRG2021-4648) and the support from Google Cloud through the Google Gemma 2 Academic Program GCP Credit Award. SA is supported by a Newton Trust Grant. MB acknowledges financial support via the Open Philanthropy Foundation funded by the Good Ventures Foundation. This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP4, project number: 276693517 and the UKRI grant: Turing AI Fellowship EP/W002981/1. MB is a member of the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.
@@ Appendix C Concept Frequency is Predictive of Performance for T2I Models
We extend the results from Fig. 3 with Figs. 11, 12, 13, 14 and 15. As with Fig. 3, due to the high concept frequency, the scaling trend is slightly less pronounced. Furthermore, we do see inconsistency in the trends for the human-rated scores retrieved from HEIM [72], hence we perform our own small scale human evaluation to check them.
@@
We assessed image-text alignment with a human study with 6 participants, where each participant was assigned 72 samples; for consistency, of the 360 total samples, we ensured 10% were assigned to 3 participants. Provided with a reference image, the participants were asked if the sample accurately depicts the prompt (see Fig. 16). Specifically, “Does the image accurately depict the above prompt?”. Three choices were provided: “Yes” (score=1.), “Somewhat” (score=0.5), and “No” (score=0.). Accuracy was computed by averaging the scores.
@@
Note on Participant Acquisition. Experiment participants, who volunteered for the study, provided informed consent. IRB approval was not obtained.
@@
Figure 16: User Interface for T2I human evaluation for text-image alignment for people concepts. See Appx. C for further details.
@@ Appendix D Concept Frequency is Predictive of Performance across Concepts only from Image and Text Domains
In all the main performance-frequency plots we have presented until now, the concept frequencies were estimated using the intersection of the image-frequencies and the text-frequencies. Here, we showcase results with using them independently in Figs. 19 and 18 respectively. We note that both independent searching methods showcase log-linear trends as before confirming our main result. We observe that the strong log-linear trend between concept frequency and zero-shot performance robustly holds across concepts derived from image and text domains independently as well.
@@
Appendix E Generalization of findings to improved VLM training objectives
We believe our main conclusions of exponential data inefficiency should hold regardless of the model architecture and the training objective for any VLM. However, to test this thoroughly, we investigated two methods that have been empirically shown to improve generalization capabilities of CLIP models: CyCLIP [49] and SLIP [87]. We use $4$ different models, each trained with either CyCLIP/SLIP on three different datasets—we then plot our main log-linear scaling results similar to Fig. 2 for CyCLIP and SLIP models—these plots are in Fig. 20. We observe for both SLIP and CyCLIP models, the log-linear scaling trends hold strong, with high Pearson correlation coefficients, further signifying the robustness of our main results. Hence, we emphasize that our main conclusions hold true even when considering multimodal models that explicitly introduce new training objectives with the aim of improving model generalization.
Figure 20: Log-linear scaling trends for SLIP and CyCLIP models
Appendix F Extended Related Work
Supplementing the discussion in the main paper, we further provide a broad overview of the surrounding literature within which our paper is mainly positioned.
Effect of Pre-training Data on Downstream Data. Several data-centric prior works [98, 47, 88, 43, 89, 76, 134, 135, 145, 117, 80, 99, 107, 108, 39, 27, 103] have highlighted the importance of pretraining data in affecting performance. Fang et al. [43] robustly demonstrated that pretraining data diversity is the key property underlying CLIP’s strong out-of-distribution generalisation behaviour. Similarly, Berlot-Attwell et al. [17] showed that attribute diversity is crucial for compositional generalization [60], namely systematicity [46]. Nguyen et al. [88] extended the Fang et al. [43] analysis to show that differences in data distributions can predictably change model performance and that this behaviour can lead to effective data mixing strategies at pretraining time. Mayilvahanan et al. [81] complemented this research direction by showing that CLIP’s performance is correlated with the similarity between training and test datasets. Udandarao et al. [129] further showed that the frequency of certain visual data-types in the LAION-2B dataset was roughly correlated to the performance of CLIP models in identifying visual data-types. McCoy et al. [82] introduced the teleological approach to understanding model generalization, showing that LLMs are more reliable on tasks and input/output instances that are more probable based on their pretraining datasets. Our findings further pinpoint that the frequency of concept occurrences is a key indicator of performance. This complements existing research in specific areas like question-answering [62] and numerical reasoning [102] in large language models, where high train-test set similarity does not fully account for observed performance levels [137]. Concurrent to our work, Parashar et al. [92] also explore the problem of long-tailed concepts in the LAION-2B dataset and how it affects performance of CLIP models, supporting our findings. In contrast to their work, we look at count separately in image and text modalities, as well as across pretraining sets, and do a number of control experiments to thoroughly test the robustness of our result. Further, our frequency estimation procedure on text captions and images independently enables us to provide a more fine-grained analysis of the pretraining data distribution like quantifying misalignment between images and texts, and assessing similarity of the different pretraining data concept distributions. Finally, our demonstration that the long tail yields a log-linear trend explicitly indicates exponential sample inefficiency in large-scale pretrained models.
Detailed Differences and Contributions compared to Parashar et al. [92]. We emphasise that we complement this prior work and point out that our work comprehensively tests the strength of the log-linear scaling trend across several datasets, spanning varying levels of data curation and dataset sizes. Further, we note that in Parashar et al. [92], the estimated frequencies are computed using only the text captions of LAION-2B. These estimated frequencies are then used as the canonical frequencies for plotting the performance-frequency curves for all the tested models (despite these models being trained on different pretraining datasets other than LAION-2B). Our work strongly showcases why this apparent asymmetry in their frequency estimation methodology should work—from Tab. 4, we show that different VLM pretraining datasets are strongly correlated in their concept distributions. Hence, in spite of Parashar et al. [92] using only LAION-2B as their source dataset for frequency estimation, their results roughly hold true because of this strong correlation across pretraining datasets. Our methodology of incorporating both images and text captions when computing the frequency estimates is crucial for explaining this. Hence, we believe that our work comprehensively generalizes and explains the findings of prior work while also providing insights into the pretraining datasets (e.g., misalignment degree and correlation of concept distributions in datasets).
Appendix G Experimental Details
G.1 Setup of Mayilvahanan et al. [81]
@@ Appendix G Experimental Details
LAION-200M is a dataset obtained by deduplicating LAION-400M by pruning exact duplicates, near duplicates, and semantically similar samples within LAION-400M [10]. The control pretraining set is created by pruning 50 million highly similar samples from LAION in the order of decreasing perceptual similarity to datapoints in ImageNet-val set. We use the 150M pretraining set for obtaining the concept distribution. We evaluate the performance of a ViT-B-32 CLIP model trained on this dataset on our downstream tasks and present our analysis on those tasks.
Appendix H Let It Wag! Test Set
H.1 Final Set of Concepts in Let It Wag!
Based on our frequency estimation pipeline from Sec. 2, we carefully curate $290$ of the least frequent concepts across LAION-400M pretraining dataset (out of the $4,029$). We then remove all the concepts that have 0 counts to ensure that our final dataset consists of concepts that have been detected atleast once in LAION-400M, this method has also been used in Kandpal et al. [62] to ensure robustness to noise in the estimation process. We then add them as our set of concepts in Let It Wag!. A few example concepts from our final list are: {Beechcraft_1900, Black_Rosy_Finch, Irish_Wolfhound, Japanese_Chin, Kentucky_Warbler, eastern_hog-nosed_snake, eel, eggnog, flatfish, isopod, kingsnake, ladle, lakeshore, letter_opener}. We release our full concept list publicly here.
High-level insights about long-tail concepts. The broad categories of the most long-tailed concepts with a few examples for each are as follows (a majority of them are also highlighted in Figs. 28, 29, 30 and 31):
  • Birds: Western Scrub Jay, Cassins Finch, Prairie Warbler, Red eyed Vireo, Veery
  • Animals: flatworm, Tibetan Mastiff, Scottish Terrier, vine snake, newt
  • Aircrafts: A300B4, A310, Falcon 900, DHC-8-300, MD-11
  • Objects: guillotine, letter opener, ladle, dust jacket
  • Plants and fungi: mexican aster, gyromitra, great masterwort, thorn apple, cape flower
  • Misc.: consomme, stratified texture, eggnog
Further statistics of Let-It-Wag!
We provide some further statistics of the test-set below.
  • Most frequent concepts: partridge (count=9489), Bank Swallow (count=9489), eel (7907)
  • Least frequent concepts: Red-necked Grebe (count=0), SR-20 aircraft (count=0), Globe-flower (count=0)
  • Median frequency of concepts: 97.5
  • Mean frequency of concepts: 1096.2
We also show the full histogram of concept frequencies for the $290$ concepts in Let-It-Wag! in Fig. 21. From the histogram, it is evident that most of the concepts in Let-It-Wag! have frequency less than $2000$. About half of the concepts in Let-It-Wag! (approx. 140) have a frequency less than $1000$. Hence, this histogram sufficiently establishes that our Let-It-Wag! dataset truly captures the long tail.
Figure 21: Histogram of concept frequencies for Let-It-Wag! Dataset
H.2 Let It Wag!: Classification Test Set Curation
@@
To ensure our “Let It Wag!” classification dataset is thoroughly cleaned and diverse, we follow a meticulous process consisting of several cleaning, filtering and verification steps:
@@
1. Diverse Sourcing: We gather images from three different online sources—Flickr, DuckDuckGo, and Bing Search—to maximize the variety of our dataset while retaining very easy-to-classify images222we use the image sourcing pipeline of C2C [96].
@@
2. Temporal Filtering: We applied a filter to only retrieve images after January 2023 to minimize overlap with images used in the pretraining datasets of multimodal models. Note this helps mitigate but does not ensure that the overlap problem is resolved.
@@ Appendix I Why and How Do We Use RAM++?
I.1 Why RAM++ and not CLIP or open-vocabulary detectors?
@@
I.2 How: Optimal RAM++ threshold for calculating concept frequencies
@@
We ablate the choice of the threshold we use for assigning concepts to images using the RAM++ model. For the given set of concepts, RAM++ provides a probability value (by taking a sigmoid over raw logits) for each concept’s existence in a particular image. To tag an image as containing a particular concept, we have to set a threshold for deciding this assignment. We test over three thresholds: {0.5, 0.6, 0.7}, showcasing quantitative and qualitative results for all thresholds in Figs. 24 and 23.
@@
We observe best frequency estimation results using the highest threshold of 0.7. This is due to the high precision afforded by this threshold, leading to us counting only the “most aligned images” per concept as hits. With lower thresholds (0.5, 0.6), we note that noisier images that do not align well with the concept can be counted as hits, leading to degraded precision and thereby poorer frequency estimation. To ensure that we do not incorrectly tag images with erroneous concepts, our primary objective is to optimize hit precision, even if it means occasionally missing out on tagging some images with correct concepts. Hence, we use 0.7 as the threshold for all our main results.
Table 5: Example GPT-4 Descriptions fed to RAM++ on a subset of downstream datasets and concepts.
Evaluation Dataset Concept GPT-4 Description
ImageNet [36] Tench A tench is a freshwater fish that typically has a greenish or brownish body with reddish fins, small scales, and a pair of barbels near its mouth. It can grow up to 70 cm long.
SUN397 [133] Alley An alley is a narrow passageway or lane between or behind buildings, which is often used for access or for parking.
UCF101 [116] Blowing_Candles A Blowing_Candles moment can be identified concisely as a moment or event where a person is blowing out the candles on a cake, typically at a birthday celebration.
Caltech101 [45] Butterfly A butterfly is a small, flying insect known for its colorful and symmetrical wings. It has a slender body, antennae and three pairs of legs.
CUB [131] Least_Auklet A Least Auklet is a small seabird with a black back and wings, white underparts, and a stubby orange bill. They also have white eye-rings and a small, rounded tail.
EuroSAT [56] Pasture Land Pasture land can be concisely identified as an open or cleared land covered with grass, clover, or the like, suitable for grazing by livestock, with little to no trees and is primarily used for agricultural purposes.
Flowers102 [90] pink primrose A pink primrose is a perennial flower featuring delicate, soft pink petals arranged around a yellow center, with bright green leaves at the base.
DTD [32] bumpy A bumpy object has an uneven or rough surface with lots of small raised areas or protuberances
Food101 [21] churro A churro is a long, thin, golden-brown pastry that is typically ridged and may be dusted with sugar.
FGVCAircraft [79] 707320 A 707320 is a model of the Boeing 707, which is a mid-size, long-range, narrow-body four-engine jet airliner.
Stanford-Cars [66] 1993 Volvo 240 Sedan The 1993 Volvo 240 Sedan can be identified by these features:1. Manufacturer: Volvo 2. Production Year: 1993 3. Model: 240 4. Body Style: 4-door sedan 5. Engine: 2.3L 4-cylinder 6. Transmission: 5-speed manual or 4-speed automatic
CIFAR100 [68] bowl A bowl is a round dish or container typically used to hold food, often deeper than a plate with a wide open top.
COCO-5K [75] metro A metro can be identified concisely as an urban railway system that operates within large cities, offering high-frequency services and utilizing multiple cars and stations.
@@
Figure 23: Qualitative Results with different RAM++ thresholds. We show qualitative examples across three different thresholds: {0.5, 0.6, 0.7} for estimating concept frequency using the RAM++ model. We note the significantly better concepts identified by the higher threshold (0.7) compared to the lower thresholds (0.5, 0.6). The images are sourced from the CC-3M dataset.
@@
Figure 24: Effect of different thresholds for determining concept frequency using RAM++. We test three different thresholds: {0.5, 0.6, 0.7} for estimating concept frequency using the RAM++ model. Note that the lower the threshold, the lower precision are the tagged concepts since lower thresholds (0.5, 0.6) lead to noisier images being counted as hits, hence reducing the hit precision for determining frequency. Despite this added noise at lower thresholds, we note that all the correlations are significantly positive across all thresholds. This further signifies the robustness of our log-linear scaling trends inspite of our frequency estimates being noisy. ** indicates that the result is significant ($p<0.05$ with two-tailed t-test.), and thus we show Pearson correlation ($\rho$) too.
I.3 GPT-4 Descriptions for each extracted concept
In addition to providing a list of concepts to the RAM++ model, we also provide a set of GPT-4 generated responses that describe each concept (please refer to Tab. 5 for examples). This ensures that we adequately cover synonyms of concepts and take into account concept hierarchies [84]. This further improves tagging precision by using visual descriptions to better identify concepts (this has been shown to enhance performance in previous works [97, 83]. We open-source these descriptions.
Appendix J Clarification regarding 00-frequency points
In all our main plots, we explicitly exclude zero-frequency concepts from our evaluations following Kandpal et al. [62], since frequency estimation is potentially noisy, leading to low recall rates (also discussed in Appx. I). However, to verify if our log-linear trends still hold when including all the zero-frequency concepts, we re-plot all our main zero-shot classification results from Fig. 2 by including the ones which have zero-frequencies— Fig. 25 showcases these results. We find our main log-linear scaling trends are retained. To further corroborate this, we present average accuracies for concepts with frequency 00 and non-zero frequency bins in Tab. 6 below. We note that average performance for the 00-frequency concepts is significantly lower than other non-zero frequency concepts, especially when compared to very high-frequency concepts. This justifies our main claim that exponentially more data is needed per concept to improve performance linearly.
Figure 25: Main Results with 00-frequency concepts. We re-plot all our main classification results from Fig. 2 by including concepts which have zero-frequencies. Note that the 00-frequency points are assimilated into the $10^{0}$ bin—in each plot, the $10^{0}$ bin (leftmost) consists of about $60-70\%$ 00-frequency concepts. We find that our main log-linear scaling trends are retained.
Table 6: Performance per frequency bin. Here, we explicitly report the average classification performance of models trained on different pretraining datasets, per frequency bin (i.e., 00-frequency concepts only, concepts with frequencies in the range ${1}{-}{10}$, ${10}{-}{100}$ etc.). We note that average performance for the 00-frequency concepts is significantly lower than other non-zero frequency concepts, especially when compared to the performance of very high-frequency concepts.
Dataset/Model Freq=0 Freq=1-10 Freq=10-100 Freq=100-1000 Freq=1000-10000
CC-3M/RN50 5.10 13.89 20.18 32.93 44.30
CC-3M/ViT-B-16 4.27 11.98 17.21 27.48 39.24
CC-12M/RN50 12.91 21.49 27.75 39.48 50.38
CC-12M/ViT-B-16 16.48 25.59 32.07 45.65 57.06
YFCC-15M/RN50 16.49 19.59 24.12 34.26 39.97
YFCC-15M/RN101 17.43 22.06 25.72 36.77 43.14
YFCC-15M/ViT-B-16 20.75 25.06 29.68 38.73 45.96
LAION-400M/ViT-B-32 47.41 46.42 50.53 55.96 65.00
LAION-400M/ViT-B-16 51.77 52.09 57.12 61.32 70.73
LAION-400M/ViT-L-14 60.44 58.87 62.43 67.63 76.65
Appendix K Misalignment Degree Results and Human Verification